Fermi National Laboratory


Accelerator Update for October 31, 2005 - November 2, 2005

Monday October 31
The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) and the Meson MCenter experiment off due to shot setup, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beam, and with NuMI and MiniBooNE taking beam.

Operations established Recycler only store 4477 at 9:28 AM with an initial luminosity of 164.17E30. This store set a New Luminosity Record.

At 9:31 AM, a Linac expert investigated a problem with a drifting quadrupole power supply (L:QT3).

At 9:37 AM, a TeV RF station (TRF1) tripped off. There was minimal effect on the colliding antiprotons.

At 11:49 AM, experimenters at CDF reported water spilling through the concrete barrier on the west side of their parking lot. A FESS expert said that there was an ICW leak in the area and that they were working on stopping the water flow.

Water Leak

FESS reported at 12:54 PM, that they had found and stopped the water leak. They began planning on how to repair the leak.

Booster experts began a 400 MeV study at 1:21 PM. This study limited all useful downstream beam.

Pbar resumed stacking with what beam there was at 2:10 PM.

Booster studies ended at 2:15 PM.

At 2:18 PM, a LCW leak was reported in Linac. A Linac technician and operators found water under a Klystron RF station (KRF1). Operations halted Linac beam while the techs began a temporary repair. They completed the repair at 3:18 PM.

Pbar resumed stacking at 3:32 PM.

MCenter resumed taking beam at 4 PM.

At 5:55 PM, a cryo system tech reported that the TeV sector E1 wet engine was failing; one of its exhaust valves was stuck closed. An expert was called in.

Operation began preparing the Recycler at 7:38 PM to stash antiprotons. The stash ended and MCenter resumed taking beam at 8:25 PM. Pbar resumed stacking five minutes later.

Tuesday November 1
The midnight shift began with Operations monitoring store 4477 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing at 12:16 AM to stash. The stash ended and Pbar resumed stacking at 1:08 AM. MCenter resumed taking beam.

At 3:07 AM, Operations reported a vacuum burst in the P1 transfer line. The vacuum recovered quickly.

The Recycler began preparing at 4:59 AM to stash. The stash ended. Pbar resumed stacking and MCenter resumed taking beam at 5:41 AM.

The Recycler began preparing at 9:32 AM to stash. The stash ended. Pbar resumed stacking and MCenter resumed taking beam at 10:24 AM.

Pbar experts began a reverse proton study at 10:31 AM.

Operations reported around 11:38 AM that the TeV sector F2 wet engine had a bearing failure.

TeV experts began an end-of-store study around 1:09 PM.

At 3:24 PM, store 4477 quenched at TeV sector F4 when Operations tried to terminate it. TeV experts investigated the quench, but couldn't find a cause.

Operations prepared CDF and DO for a collision hall access at 3:39 PM.

Pbar reverse proton studies ended at 4:31 PM.

At 4:40 PM, a Booster RF station (BRF15) tripped off for the first of six times in about forty-five minutes. It reset each time.

At 4:44 PM, a TeV RF expert reported that the TRF1 station had a bad Power Amplifier (PA). The expert had the PA changed out two hours later.

At 4:46 PM, Operations reported that they couldn't get a Pbar quadrupole power supply (Q109V) to turn on. An expert managed to get Q109 up and running on one power supply, but Q109V remained off.

Pbar resumed stacking at 5:35 PM.

Techs completed the TeV F2 wet engine work at 9:33 PM.

Operations began the TeV turn on at 10:36 PM.

Operations put the TeV through a dry squeeze at 11:36 PM.

Wednesday November 2
The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations began putting the TeV through a wet squeeze at 12:06 AM.

Operations put the TeV into shot setup at 12:48 AM.

At 1:04 AM, the MiniBooNE Horn tripped off. An expert came in an investigated.

At 2:32 AM, the NuMI chiller tripped off. An expert directed Operations to reset it.

At 3:11 AM, the NuMI chiller tripped off again. The expert told Operations to reset it again.

Operations established store 4480 at 3:18 AM with an initial luminosity of 146.74E30.

The Recycler began preparing at 3:28 AM to stash. The stash was completed and Pbar and MCenter resumed normal operations by 4:13 AM.

The Recycler began preparing at 7:34 AM to stash.

The Plans for Wednesday
The plans for today are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/2/2005   email Fermilab